# Whole-Body Model-Predictive Control of Legged Robots with MuJoCo

### v1

@@ Abstract
We demonstrate the surprising real-world effectiveness of a very simple approach to whole-body model-predictive control (MPC) of quadruped and humanoid robots: the iterative LQR (iLQR) algorithm with MuJoCo dynamics and finite-difference approximated derivatives. Building upon the previous success of model-based behavior synthesis and control of locomotion and manipulation tasks with MuJoCo in simulation, we show that these policies can easily generalize to the real world with few sim-to-real considerations. Our baseline method achieves real-time whole-body MPC on a variety of hardware experiments, including dynamic quadruped locomotion, quadruped walking on two legs, and full-sized humanoid bipedal locomotion. We hope this easy-to-reproduce hardware baseline lowers the barrier to entry for real-world whole-body MPC research and contributes to accelerating research velocity in the community. Our code and experiment videos will be available online at: https://johnzhang3.github.io/mujoco_ilqr.
@@ I Introduction
This paper aims to reduce this gap by providing an open-sourced baseline MPC algorithm and real-world legged robot implementation built on the MuJoCo physics engine [6], a standard, easy-to-use open-source robotics simulator. We show that a standard gradient-based MPC algorithm, in particular the iterative LQR (iLQR) algorithm, based on MuJoCo is surprisingly capable of solving a variety of challenging real-world tasks such as bipedal locomotion on a quadruped and full-sized humanoid robot in real time. We hope that this effort lowers the barrier to entry for further model-based control research on legged robot hardware and eventually leads to accelerated research momentum.
@@ II-A Iterative Linear-Quadratic Regulator
where and are the current solution, is the feedback gain matrix at time index , is an improvement to the current control, and is the line search step size. This single-shooting formulation, which only optimizes over the controls and recovers the corresponding states by integrating over the discrete-time dynamics . and are the running and terminal costs, respectively. In each iteration, derivatives of cost and dynamics w.r.t to the control sequence are computed around the current solution points in a process called linearization to form a subproblem with quadratic cost and linear constraints. The term iterative LQR [17, 18], or iLQR, generally refers to the Gauss-Newton approximations of the original DDP algorithm that is often more computationally efficient. The iLQR algorithm can also be modified to handle control limits [19], state constraints [13, 15], and contact [20].
@@ II-C Open-Source Tooling for Model-Based Control
Outside the MuJoCo ecosystem, Pinocchio [25] has become a popular toolbox in the community for efficiently computing rigid body dynamics and its derivatives. This dynamics toolbox has also enabled a variety of nonlinear trajectory optimization libraries [33, 34] designed for contact-rich robotics tasks like legged locomotion. Additionally, Drake [5] provides modeling features that solve accurate contact and friction dynamics at the expense of high computation costs for model-based control and verification.
@@ III-A MuJoCo Soft Contact Model and Derivatives
To approximate the model derivatives, we use the forward difference method because it requires only one additional simulation evaluation per dimension compared to two in centered difference , where is an arbitrary function with input and is the finite different tolerance.
III-B MuJoCo iLQR
We build on the open-source MuJoCo MPC [30] toolbox which implements the iLQR algorithm described in Sec. II-A and deploy on Unitree quadruped and humanoid robots, Fig. 2. Due to the nature of online motion planning, we solve a single iLQR iteration before returning the current-best nominal trajectories and feedback policy without convergence checks. The solution is then used to warm start the next iteration with new state information from the robot. Additional features such as regularization and parallel linesearch are also implemented.
@@ IV-A Contact Modeling
We address this issue by increasing the impratio, which roughly corresponds to the solver’s ability to trade off sliding versus penetration, value from the default to . This change also has the added effect of increasing the solve times of the simulation problem and, as a result, the iLQR iteration times. For our quadruped system, the time for a single iLQR iteration increases from ms to ms on a th-gen Intel i7 CPU. We do not find the added compute time to be an issue during real-world deployment. Interestingly, the moderate ground penetration from the softness of the contact model does not cause issues in the sim-to-real transfer in our experience.
IV-B Actuator Impedance
MuJoCo simulation allows easy definition of the joint-level impedance in well-defined units. However, to our knowledge, the impedance parameters on the real-world robots are provided in black-box units, which results in a non-trivial model mismatch. We resolve this for each robot with manual tuning, but more systematic methods for identifying joint-level system dynamics will be critical for maximizing robot capabilities without risking damage.
@@ V Experiments and Results
This section presents the interactive GUI setup on robot hardware platforms and a variety of hardware experiments. We start by demonstrating our system on basic quadrupedal locomotion on Go1 and Go2 robots in Sec. V-B. Next, we show that MuJoCo iLQR naturally extends to open-loop unstable tasks like quadruped walking on two legs in Sec. V-C. Finally, we deploy our system on a human-sized Unitree H1 humanoid robot in Sec. V-D.
@@ V-A Interactive GUI for Real-World Legged Robots
We interface with the MuJoCo MPC GUI and iLQR planner via Python for robot hardware deployment. The interface takes the latest state estimation and receives the current control. The state estimation is computed by fusing Optitrack MoCap [38] position and attitude measurements at Hz via ROS and the robot joint angle position and velocity measurements at Hz. The floating base linear and angular velocities in the body frame are calculated by applying a low-pass filter to finite difference values of position and attitude measurements. The planner’s actions are communicated to the robot via Unitree SDK for the Go1 robot and Unitree SDK 2 for the Go2 and H1 robots. All robot controls are represented as joint targets and tracked using Unitree internal low-level PD controllers. We compute the MPC policies on a desktop equipped with a th generation Intel i CPU chip with cores, a different CPU from earlier in the paper. We update the iLQR policy at Hz and use the TV-LQR policy to stabilize the robots between solves at Hz. Unless otherwise noted, we use a default prediction horizon of s and discretize the dynamics at Hz in the iLQR planner.
@@ ACKNOWLEDGMENT
The authors thank Arun Bishop, Swaminathan Gurumurthy, and Shuo Yang for insightful discussions and feedback. This work is supported by a Google grant.

### v2

@@ Abstract
We demonstrate the surprising real-world effectiveness of a very simple approach to whole-body model-predictive control (MPC) of quadruped and humanoid robots: the iterative linear-quadratic regulator (iLQR) algorithm with MuJoCo dynamics and finite-difference approximated derivatives. Building upon the previous success of model-based behavior synthesis and control of locomotion and manipulation tasks with MuJoCo in simulation, we show that these policies can easily generalize to the real world with few sim-to-real considerations. Our baseline method achieves real-time MPC while leveraging whole-body dynamics and collision detection on a variety of hardware experiments, including dynamic quadruped locomotion, a quadruped walking on two legs, and full-sized humanoid bipedal locomotion. Additionally, our GUI system enables users to interactively update robot behavior in real-time on the robot hardware, making task-specific parameter tuning easy and intuitive. Our code and experiment videos are available online at: https://johnzhang3.github.io/mujoco_ilqr.
@@ I Introduction
This paper aims to reduce this gap by providing an open-sourced baseline MPC algorithm and real-world legged robot implementation built on the MuJoCo physics engine [6], a standard, easy-to-use open-source robotics simulator. We show that a standard gradient-based MPC algorithm, in particular the iterative LQR (iLQR) algorithm, based on MuJoCo is surprisingly capable of solving a variety of challenging real-world tasks such as bipedal locomotion on a quadruped and full-sized humanoid robot in real time. By leveraging the efficient C implementation of MuJoCo as the backend for the forward model and derivative computations, our real-time MPC approach reasons about both whole-body dynamics and collision detection in the model, something difficult to achieve in previous open-source whole-body MPC algorithms [16, 17]. Additionally, we design an interactive GUI system that enables users to quickly modify key MPC parameters and observe real-world robot behaviors alongside a simulated twin. We hope that this effort lowers the barrier to entry for further model-based control research on legged robot hardware and eventually leads to accelerated research momentum.
@@ II-A Iterative Linear-Quadratic Regulator
where and are the current solution, is the feedback gain matrix at time index , is an improvement to the current nominal control, and is the line search step size. This single-shooting formulation, which only optimizes over the controls and recovers the corresponding states by integrating over the discrete-time dynamics . and are the running and terminal costs, respectively. In each iteration, derivatives of cost and dynamics w.r.t to the control sequence are computed around the current solution points in a process called linearization to form a subproblem with quadratic cost and linear constraints. The term iterative LQR [19, 20], or iLQR, generally refers to the Gauss-Newton approximations of the original DDP algorithm that is often more computationally efficient. The iLQR algorithm can also be modified to handle control limits [21], state constraints [13, 15], and contact [22].
@@ II-C Open-Source Tooling for Model-Based Control
Outside the MuJoCo ecosystem, Pinocchio [27] has become a popular toolbox in the community for efficiently computing rigid body dynamics and its derivatives. This dynamics toolbox has also enabled a variety of nonlinear trajectory optimization libraries designed for contact-rich robotics tasks like legged locomotion. For example, OCS2 [34] offers open-source software for quadruped locomotion with fixed contact modes, while Crocoddyle [16] and Aligator [35] offer more versatile control for quadruped and humanoid robots. However, these works do not leverage off-the-shelf simulators widely used in the robotics community and have a steep learning curve, which limits their reach and impact. On the other hand, Drake [5] provides advanced modeling and simulation features for accurate contact and friction dynamics for verification is not fast enough for real-time control.
In comparison to prior projects (Table I), our approach bridges the current gap in tooling for model-based robotic control by enabling general controllers for both quadruped and humanoid robots using a popular, off-the-shelf, and fast robotics simulator. An additional advantage for using a mature simulator is the readily available collision detection algorithms that we can leverage during contact-rich planning and control. Additionally, we provide an interactive GUI for real-time control that enables rapid developments of robot behaviors in the real world.
@@ III-A MuJoCo Soft Contact Model and Derivatives
To approximate the model derivatives, we use the forward difference method
as it requires only one additional simulation evaluation per dimension compared to two in centered difference, where is an arbitrary function with input and is the finite different tolerance.
III-B Derivative Computation with MuJoCo with MuJoCo
To solve a single iteration iLQR problem from Eq. 1, we take a second-order Taylor expansion and solve the resulting subproblem via Dynamics Programming [18, 20].
We define the cost function as the following:
where is a residual vector to be reduced when solving the problem, is the norm function that returns a non-negative scaler, and is non-negative scaler weight defining the importance of a residual term.
We compute the exact cost gradients:
and approximate the cost hessians as:
where the derivatives of the norm n (i.e. , ) are computed analytically and the Jacobians of the residual r (i.e. , ) are computed via finite difference, Eq. 3.
We take the original nonlinear discrete-time dynamics:
and linearize as:
where dynamics Jacobians and are computed via finite difference, Eq. 3. Note the because residuals in Eq. 4 are implemented as MuJoCo sensors, we can efficiently compute all the Jacobians , , , and via a single call to the MuJoCo finite difference utilities function [6].
Once all derivatives are computed, we solve the resulting problem using the Riccati-recursion that produces the updated nominal controls and a time-varying linear feedback policy , Eq. 2. We perform this update once before returning the current-best nominal trajectories and feedback policy without convergence checks. We then use the previous solution to warm start a new iLQR iteration with the latest state estimation, Fig. 2. Additionally, our interactive GUI allows the user to update the residual terms such as target height, goal positions, etc and adjust the weights assigned to each residual term in real-time on the robot, Fig. 3.
@@ IV-A Contact Modeling
We address this issue by increasing the impratio value, which roughly corresponds to the solver’s ability to trade off sliding versus penetration, from the default to . This change also has the added effect of increasing the solve times of the simulation problem and, as a result, the iLQR iteration times. For our quadruped system, the time for a single iLQR iteration increases from ms to ms on a th-gen Intel i7 CPU. We do not find the added compute time to be an issue during real-world deployment. Interestingly, the moderate ground penetration from the softness of the contact model does not cause issues in the sim-to-real transfer in our experience.
IV-B State and Action Representation
We represent the state of the robot in the standard way, with a floating base position and attitude quaternion, followed by joint angles and corresponding linear, angular, and joint-angle velocities. We include a low-level joint-space PD controller in the dynamics model of the robot such that the inputs to the model that are optimized by iLQR are joint angle references. As a result, our approach achieves direct whole-body control of the robot without the model hierarchies commonly seen in traditional model-based MPC algorithms [24]. By leveraging a fast off-the-shelf simulator, our method is simpler and more accessible than methods relying reduced-order models and hierarchical control approaches.
@@ V Experiments and Results
This section presents the interactive GUI setup on robot hardware platforms and a variety of hardware experiments. We start by demonstrating our system on basic quadrupedal locomotion on Go1 and Go2 robots in Sec. V-B. Next, we show that MuJoCo iLQR naturally extends to open-loop unstable tasks like quadruped walking on two legs in Sec. V-C. Finally, we deploy our system on a human-sized Unitree H1 humanoid robot in Sec. V-D. Our open-source software and experiment videos are available at:
Ommitted for Anoynymous Review
@@ V-A Interactive GUI for Real-World Legged Robots
We interface with the MuJoCo MPC GUI and iLQR planner via Python for robot hardware deployment. The interface takes the latest state estimation and receives the current control. The state estimation is computed by fusing Optitrack MoCap [39] position and attitude measurements at Hz via ROS and the robot joint angle position and velocity measurements at Hz. The floating base linear and angular velocities in the body frame are calculated by applying a low-pass filter to finite difference values of position and attitude measurements. The planner’s actions are communicated to the robot via Unitree SDK for the Go1 robot and Unitree SDK 2 for the Go2 and H1 robots. All robot controls are represented as joint targets and tracked using Unitree internal low-level PD controllers. We compute the MPC policies on a desktop equipped with a th generation Intel i CPU chip with cores, a different CPU from earlier in the paper. We update the iLQR policy at Hz and use the TV-LQR policy to stabilize the robots between solves at Hz. Unless otherwise noted, we use a default prediction horizon of s and discretize the dynamics at Hz in the iLQR planner.Note that, while we use a desktop-level CPU in our experiments, we see similar policy update frequencies from recent top-end laptop Arm-based CPUs such as the M-series chips from Apple, but leave deployment for future work.
@@ V-B Quadruped Locomotion
The residual terms for the locomotion task include tracking position and orientation while keeping balance and maintaining a desired torso height. Control efforts are penalized to minimize energy usage. We also include a nominal gait pattern in the residual, but note that this is not implemented as a constraint. As a result, the iLQR solver is free to discover new contact modes if they reduce the overall cost. More detailed descriptions of the residual terms can be found in Tab. II. Bipedal locomotion and humanoid tasks follow a similar structure.
@@ V-C Bipedal Locomotion for Quadruped Robots
| Term | Description |
| Upright | Keeps torso orientation upright (z-axis pointing up) |
| Height | Controls torso height relative to average foot position |
| Position | Tracks head position to target location (x, y, z) |
| Gait | Controls foot lifting patterns during gait cycles (one per foot) |
| Balance | Keeps capture point within support polygon (x, y components) |
| Effort | Penalizes actuator forces to minimize energy consumption |
| Posture | Keeps joints near home configuration |
| Yaw | Controls heading direction (x, y components of heading vector) |
| Angular | Controls rotational dynamics (x, y, z components) |
@@ ACKNOWLEDGMENT
Omitted for Anonymous Review